OpenStudy (anonymous):

One cubic centimeter (1.0 cm3) of water has a mass of 1.0 10-3 kg. (a) Determine the mass of 1 m3 of water. Answer: 1000 kg (b) Assuming biological substances are 98% water, estimate the mass of a (spherical) cell with a diameter of 1.0 µm. ___ kg (c) Assuming biological substances are 98% water, estimate the mass of a human kidney. Take a kidney to be roughly a sphere with a radius of 4 cm. ____ kg (d) Assuming biological substances are 98% water, estimate the mass of a fly. Take a fly to be roughly a cylinder 4 mm long and 2 mm in diameter. ___ kg

OpenStudy (anonymous):

\( \Large 1.0 µm=0.0001 cm\) Volume of a cell: \(\Large V=\frac{4}{3}\pi (\frac{d}{2})^{3}=\frac{4}{3}\pi (\frac{0.0001}{2})^{3} \quad (cm^{3})\) Because Assuming biological substances are 98% water,so the mass of a cell is: \(\large m=\frac{98}{100}\times V\times 1.0\times10^{-3}=\frac{98}{100}\times \frac{4}{3}\pi (\frac{0.0001}{2})^{3}\times 1.0\times10^{-3} \approx 5.13(kg)\)

OpenStudy (anonymous):

\(C.\) \(\Large 0.26 kg\) \(\Large d. V=\pi (\frac{d}{2})^{2}h=\pi (\frac{0.2}{2})^{2}0.4\) The mass of a fly: \(\Large m=\frac{98}{100}\times V\times1.0\times10^{-3}=1.23 (kg)\)

OpenStudy (vincent-lyon.fr):

A 1.23 kg fly is not possible!

OpenStudy (anonymous):

Sorry, I got mistake. The mass of a cell is: \(\Large m=5.13\times10^{-18}\) And,The mass of a fly: \(\Large m=1.23\times10^{-5} (Kg)\)